OpenStudy (anonymous):

Medal A personal trainer has a supply of weight plates to make custom barbells to use for weight-training. The trainer adds 13 3/4 pounds of plates to each side of the dumbbell bar, which itself weighs 7 1/2 pounds. Write a ratio comparing the weight added to each side of the dumbbell bar to the weight of the bar itself, using a fraction in simplest terms.

OpenStudy (anonymous):

anyone can you help me

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

its 9.75

zepdrix (zepdrix):

Yah yah yah I misread the question :( He added `13 and 3/4` pounds of plates to each side. So that's supposed to be a mixed number.

zepdrix (zepdrix):

\[\Large\rm \frac{\text{weight added to each side}}{\text{weight of the bar}}=\frac{13\frac{3}{4}}{7\frac{1}{2}}\]So we want our ratio to look something like this, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

They want it in simplest terms though, We would like to rewrite our `mixed numbers` as improper fractions.

OpenStudy (anonymous):

so its 55/4 15/2

zepdrix (zepdrix):

\[\Large\rm \frac{55/4}{15/2}\]Ok great! Do you remember how to deal with fractions when they're being divided? Did you learn `keep change flip` or something similar maybe?

OpenStudy (anonymous):

yes we do that

zepdrix (zepdrix):

Mmm k good :) So you `keep` the 55/4 the same, you `change` the operation from division to multiplication, and `flip` the bottom fraction.

zepdrix (zepdrix):

\[\Large\rm \frac{55}{4}\times \frac{2}{15}\]Like ummm, like that, yah?

OpenStudy (anonymous):

yes i understand

zepdrix (zepdrix):

You COULD multiply the tops and bottoms separately, but it's maybe a good idea to simplify things BEFORE multiplying.

zepdrix (zepdrix):

\[\Large\rm \frac{55}{\cancel{4}2}\times \frac{\cancel{2}1}{15}\]Do you understand what I did right there? Dividing a 2 out of the top and bottom.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i understand

zepdrix (zepdrix):

How bout with the numbers ending in a 5, can we do something similar with those? What do you think? :) Do it!!

OpenStudy (anonymous):

divide them by 5

zepdrix (zepdrix):

Yes yes yes. So do it :3 What do you get for an answer?

OpenStudy (anonymous):

11/3

OpenStudy (anonymous):

am i right?

zepdrix (zepdrix):

\[\Large\rm \frac{55}{\cancel{4}2}\times \frac{\cancel{2}1}{15}\] \[\Large\rm \frac{55}{2}\times \frac{1}{15}\] \[\Large\rm \frac{\cancel{55}11}{2}\times \frac{1}{\cancel{15}3}\] \[\Large\rm \frac{11}{2}\times \frac{1}{3}\]Ok good, taking a 5 out gives us that. Finish up the multiplication though ^^

OpenStudy (anonymous):

so its 11/6

zepdrix (zepdrix):

Yayyy good job \c:/

OpenStudy (anonymous):

is this the ratio

zepdrix (zepdrix):

`Write a ratio` ... `using a fraction in simplest terms.` Yes! We've finally got it in the simplest form! :)

OpenStudy (anonymous):

ok thank you :)